# ETTRL: Balancing Exploration and Exploitation in LLM Test-Time Reinforcement Learning Via Entropy Mechanism

### v1

@@ 3.3 EAR: the exploration and exploitation of Reward learning
During TTRL training, the policy model employs parallel sampling and estimates pseudo-labels via majority voting to generate feedback rewards. In the initial training phase, taking the AIME task as an example, the majority ratio remains below 10% (as illustrated in Figure 3), indicating that only a small fraction of samples obtain positive rewards. After in-group advantage normalization, these limited samples unexpectedly achieve disproportionately high advantages.
@@
Cui et al. (2025) investigated the impact of entropy mechanisms on reinforcement learning, noting that response entropy can serve as a metric for assessing a model’s confidence in its outputs. Building on this work, we introduce an entropy-based mechanism to quantify the model’s confidence in its responses, aiming to mitigate the overconfidence phenomenon observed during early training stage. Our objective is to attenuate the update magnitude for samples that obtain high reward advantages yet exhibit insufficient intrinsic information.
However, the absolute value of entropy is not a standardized metric, as its magnitude is influenced by multiple factors including the datasets, sampling parameters, and policy models. To quantify the relative uncertainty of a single response, we use the average response entropy within the group as a reference baseline. Responses with entropy exceeding this average are considered to exhibit relatively low confidence. We observe that under low majority ratios, certain samples attain a positive advantage via majority voting despite high response entropy, indicating that consensus is reached with low model confidence. To suppress such positive reward effects arising from low-confidence scenarios, we define an advantage scaling coefficient based on relative entropy, as formalized in Equation 12.
@@
$\hat{A}^{res}_{i,t}=\text{clip}\bigl{(}Y_{i},1-\alpha,1+\beta\bigr{)}*\hat{A}_{i,t}$ (14)
Here, $H_{resp}$ denotes the response entropy (defined in Equation 5), and $G$ represents the number of rollouts. To balance exploration and exploitation, the advantage is amplified when response entropy exceeds the average, thereby encouraging greater exploratory behavior. Conversely, when the entropy falls below the average, it proportionally reduces the advantage, thereby diminishing the magnitude of policy updates for uncertain responses. Moreover, to ensure training stability, we apply clipping to the scaling factor as formalized in Equation 12. Leveraging the standardized response entropy as an explicit conditioning signal, we introduce Adv-Res method — an advantage estimation algorithm that modulates the scaling intensity of advantage values via relative-entropy-based shaping.
Regarding the constraint of advantages in reinforcement learning, scaling is not the sole approach; clipping is another widely adopted technique. Consequently, we conduct comparative experiments on advantage clipping (Adv-Clip), which constrains advantage values within a predefined range. The formulation of Adv-Clip is presented below:

### v2

@@ 3.3 EAR: the exploration and exploitation of Reward learning
During TTRL training, the policy model generates pseudo-labels via majority voting over sampled responses. In the initial phase, however, the majority ratio is often extremely low (e.g., below 10% on AIME), meaning that only a small portion of samples obtain positive rewards. After normalization within each rollout group, these few “lucky” samples are assigned disproportionately large advantages, which in turn amplify their gradients.
@@
To counteract this instability, we adopt Adv-Clip as the primary regularization strategy. The core idea of Adv-Clip is straightforward yet highly effective: it constrains the magnitude of the advantage values within a predefined range, thereby directly suppressing extreme updates in the early stages of training. Formally, the clipped advantage is expressed as:
By bounding the scale of policy gradients, Adv-Clip prevents a small number of noisy or low-confidence samples from dominating optimization. This mechanism is particularly crucial in the early phase, when pseudo-label accuracy is low and unstable. Empirically, we observe that clipping stabilizes learning curves, reduces the risk of divergence, and maintains sufficient exploration capacity for later stages. Conceptually, Adv-Clip acts as a safeguard, ensuring that the model does not prematurely collapse its exploration due to overconfident yet unreliable reward signals.
While clipping effectively mitigates overconfidence, it does not exploit finer-grained information about the reliability of each response. Cui et al. (2025) investigated the impact of entropy mechanisms on reinforcement learning, noting that response entropy can serve as a metric for assessing a model’s confidence in its outputs.
To further refine advantage estimation, we introduce an entropy-based mechanism Adv-Res as a complementary strategy, which is expressed as:
$\hat{A}^{res}_{i,t}=Y_{i}*\hat{A}_{i,t}$ (13)
@@
Here, $H_{resp}$ denotes the response entropy (defined in Equation 5), and $G$ represents the number of rollouts. Adv-Res leverages response entropy to assess relative confidence: responses with higher-than-average entropy are considered uncertain, and their advantages are down-weighted, while low-entropy responses receive slightly amplified updates. This soft adjustment enriches the exploration–exploitation balance and yields additional improvements in performance.